OpenStudy (science_ally):

How do I combine like terms in parenthesis ? Solve the equations by combining like terms. Show all steps. 2(2x-4)+x=7 and also 8=3(3x+8)-x

OpenStudy (science_ally):

@akonkel

OpenStudy (anonymous):

2(2x-4)+x=7 you use the distributive property... 2*2=4 2*4=8 so you now have... 4x-8+x=7 then you add 4x+x and you get.. 5x so now its 5x-8=7 to solve you would add 8 to both sides... 5x-8(=8)=7(+8) which is... 5x=15 divide both sides by 5... x=3

OpenStudy (akonkel):

first u want to do whats in the ( ) first remember PEMADS.....

OpenStudy (akonkel):

woop never mind someone got it already :/ ^-^

OpenStudy (science_ally):

thnx

OpenStudy (anonymous):

Do you need help with the second one or do you got it?

OpenStudy (science_ally):

yes I'm doing more on paper now and trying to create my own 2 problems I appreciate it soo much.

OpenStudy (anonymous):

Wait lol was that a yes to u need help with it?

OpenStudy (science_ally):

yes please

OpenStudy (anonymous):

Oh okay one second.

OpenStudy (anonymous):

3(3x+8)-x= 8 distribute the 3 to both numbers 3*3= 9 3*8= 24 9x+24-x=8 now do 9x-x= 8x 8x+24=8 subtract 24 from both sides 8x+24(-24)=8(-24) 8x=-16 divide both sides by 8.. x=-2

OpenStudy (science_ally):

thnk u so much : ) can you tell me if this makes sense : 16a+4a+4=(-64) ?

OpenStudy (anonymous):

I just did it on paper and your answer would be -17/5 i guess it makes sense if you get an answer

OpenStudy (science_ally):

thnx

OpenStudy (anonymous):

No problem :)